OpenStudy (anonymous):

Find the y-intercept of the straight line passing through the point P = (5, 3) and perpendicular to y + 5x = 1 . a. y-intercept = −2 b. y-intercept = −4 c. y-intercept = 2 d. y-intercept = 4 e. y-intercept = 9/5

OpenStudy (anonymous):

Perpendicular means Product of their Slopes = -1 ie m1*m2=-1

OpenStudy (anonymous):

y + 5x = 1 y=-5x + 1 slope = -5 -5 * m2 = -1 m2 = 1/5

OpenStudy (anonymous):

Nw use Point Slope form....

OpenStudy (anonymous):

y=-5x+1/5?

OpenStudy (anonymous):

y-3=1/5 ( x-5) y-3 = x/5-1 y=x/5 + 2 so...y intercept = 2